OpenStudy (anonymous):

How do you factor 10a^2+a-3

OpenStudy (zzr0ck3r):

10a^2+a-3 = 10a^2-5a+6a-3 = 5a(2a-1)+3(2a-1) = (5a+3)(2a-1)

OpenStudy (zzr0ck3r):

make sence? I used the "AC" method.

OpenStudy (anonymous):

can you tell me by sum and product?

OpenStudy (zzr0ck3r):

i dont know what that means

OpenStudy (anonymous):

okay nvm, thank you though :)

OpenStudy (zzr0ck3r):

ok I read a little sum of roots is -b/a so in our situation we have -1/10 sum of roots says the two roots we find should add to -1/10, the roots i found are 5a+3=0 so a = -3/5 2a-1 = 0 so a = 1/2 so we add them-3/5+1/2 = -6/10 + 5/10 = -1/10 so sum of roots works out product of roots says that (-3/5)(1/2) should equal c/a or -3/10 so we check (-3/5) * (1/2) = -3/10

OpenStudy (zzr0ck3r):

im not sure if this is what your asking, but hope it helps